# Characterizing Density and Gravitational Potential Fluctuations of the Interstellar Medium

### v1

@@ Appendix B Fitting a Fluctuation Vertical Profile
As detailed in Equation 39, the separable density model we apply throughout Section 4 assumes that the vertical profile $\zeta$ of the $k=0$ component of the density also applies to the fluctuations, while in reality, each Fourier component really possesses its own vertical profile $\zeta_{\mathbf{k}}$ that need not be equal to the mean $\zeta_{k=0}$. While the simple “single-$\zeta$” model we present in the main text corresponds reasonably well with what is measured in the simulations, below we describe a simple extension of the model allowing for the $k=0$ and $k>0$ components of the density to have distinct vertical profiles, which we will denote $\zeta$ and $\zeta^{>}$ respectively. Of course, further subdividing $\zeta^{>}$ into several individual $\zeta_{\mathbf{k}}$ would improve the fit, but at the cost of significantly complicating the model, so we refrain from doing so here.
@@
$\delta\rho_{\mathrm{m}}(\mathbf{x},z,t)=\frac{\overline{\Sigma}(t)}{2h_{\mathrm{m}}^{>}(t)}\zeta_{\mathrm{m}}^{>}(\tilde{z})\delta(\mathbf{x},t),$ (B1)
@@
$\zeta_{\mathrm{m}}^{>}(\tilde{z})=(1-w^{>})\mathrm{sech}^{2}\left(\frac{\tilde{z}}{\alpha_{\mathrm{s}}^{>}}\right)+w^{>}e^{-|\tilde{z}|/\alpha_{\mathrm{e}}^{>}},$ (B2)
but this time to the fluctuation power spectra, i.e., we fit a model of the form Equation 54 to the measured potential power spectrum $P_{\delta\phi}(\tilde{k},\tilde{z})$, except using $\zeta_{\mathrm{m}}^{>}$ in place of $\zeta_{\mathrm{m}}$ in e.g., Equation 48.
@@
We summarize the results of this fitting process in Table 6, and compare the fitted function $\zeta_{\mathrm{m}}^{>}$ to the model $\zeta_{\mathrm{m}}$ derived from the $k=0$ density component, as well as the mean vertical profile $\zeta$ as measured from the simulation in Figure 20. We see that the fluctuations are much more strongly concentrated around the midplane, as evident from the results of Section 4.2 and anticipated physically (since the feedback driving the fluctuations is strongest at the midplane).
@@
In Figure 21 and Figure 22 we compare the potential fluctuation power spectrum $P_{\delta\phi}$ measured in the simulation with both the $\zeta_{\mathrm{m}}$ and $\zeta_{\mathrm{m}}^{>}$ vertical profile models. In each figure, the mean vertical profile $\zeta_{\mathrm{m}}$ is shown as a black dashed curve as in the main text, and the fitted $\zeta_{\mathrm{m}}^{>}$ is shown as a light blue dashed curve. In Figure 21, we see that the fits are both consistent with the measurements to within ±$\pm 1\sigma$, but highlight that fitting for the $k>0$ vertical structure directly allows for a much closer match to the midplane amplitude of the fluctuations — for instance, the blue dashed curves are much closer to the simulation curves in panels (c) and (d). However, the $k>0$ fit is somewhat less accurate than the $k=0$ vertical structure fit at larger $|\tilde{z}|$. In Figure 22, we again see that at the midplane (top curve in each panel), the $\zeta_{\mathrm{m}}^{>}$ fit is closer to the simulation results, particularly at the smallest $k$ values that contain most of the power, although once again the fits suffer at larger $|\tilde{z}|$ (the lower two curves in each panel).

### v2

@@ Appendix C Fitting a Fluctuation Vertical Profile
As detailed in Equation 39, the separable density model we apply throughout Section IV assumes that the vertical profile $\zeta$ of the $k=0$ component of the density also applies to the fluctuations, while in reality, each Fourier component really possesses its own vertical profile $\zeta_{\mathbf{k}}$ that need not be equal to the mean $\zeta_{k=0}$. While the simple “single-$\zeta$” model we present in the main text corresponds reasonably well with what is measured in the simulations, below we describe a simple extension of the model allowing for the $k=0$ and $k>0$ components of the density to have distinct vertical profiles, which we will denote $\zeta$ and $\zeta^{>}$ respectively. Of course, further subdividing $\zeta^{>}$ into several individual $\zeta_{\mathbf{k}}$ would improve the fit, but at the cost of significantly complicating the model, so we refrain from doing so here.
@@
$\delta\rho_{\mathrm{m}}(\mathbf{x},z,t)=\frac{\overline{\Sigma}(t)}{2h_{\mathrm{m}}^{>}(t)}\zeta_{\mathrm{m}}^{>}(\tilde{z})\delta(\mathbf{x},t),$ (C1)
@@
$\zeta_{\mathrm{m}}^{>}(\tilde{z})=(1-w^{>})\mathrm{sech}^{2}\left(\frac{\tilde{z}}{\alpha_{\mathrm{s}}^{>}}\right)+w^{>}e^{-|\tilde{z}|/\alpha_{\mathrm{e}}^{>}},$ (C2)
[…]
@@
We summarize the results of this fitting process in Table 6, and compare the fitted function $\zeta_{\mathrm{m}}^{>}$ to the model $\zeta_{\mathrm{m}}$ derived from the $k=0$ density component, as well as the mean vertical profile $\zeta$ as measured from the simulation in Figure 20. We see that the fluctuations are much more strongly concentrated around the midplane, as evident from the results of Section IV.2 and anticipated physically (since the feedback driving the fluctuations is strongest at the midplane).
@@
[…]